Fuente, Inc., has identified an investment project with the following cash flows.
Year Cash Flow
1 $675             
2   900             
3 1,075             
4     1,425             
a.

If the discount rate is 9 percent, what is the future value of these cash flows in year 4?

      

b.

What is the future value at a discount rate of 19 percent?

      

c.

What is the future value at discount rate of 29 percent?
Answer #1

D6 for =SUM(D2:05) 1 year FV@9% $ Cash flows 675.00 $ 900.00 1,075.00 $ 1,425.00 Future value 1.2950 $ 874.14 1.1881 $ 1,069.

D6 fix =SUM(D2:05) 1 year 21 32 4 3 5 4 Cash flows 675 900 1075 1425 FV@9% =C3*1.09 =C4*1.09 =C5*1.09 Future value =B2*C2 =B3